Twice a certain natural number exceeds eight times its reciprocal by fifteen. What is the number?
I know that a number can be represented by x and twice that number is 2x but I do not know how to represent a reciprocal in a word equation.

Found 2 solutions by solver91311, edjones:
Answer by solver91311(16877)





Multiply both sides by , collect terms on the left, and factor:







Finish solving for , and then discard the negative root (you are looking for a natural number if you remember)


John

My calculator said it, I believe it, that settles it


Answer by edjones(7569)

Let the number be x.
2x=8/x + 15
2x^2=8+15x Multiply each side by x.
2x^2-15x-8=0
2x^2+x-16x-8=0
x(2x+1)-8(2x+1)=0
(x-8)(2x+1)=0
x=8
.
Ed